solve please
-36x/24x=

Answer by jim_thompson5910(35256)

-36x/24x = (-36/24)(x/x) = (-3/2)(1) = -3/2


So -36x/24x = -3/2


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim